OpenStudy (anonymous):

If you were asked to convert 25 kg to the unit mg, which of the following would be the first fraction used in the conversion? Options include 1kg/10^3.... 10^3g/1kg.... 10^3kg/1mg.... & 10^3mg/1kg

OpenStudy (anonymous):

I don't think any of these are correct. There are 1 x 10^3 mg in a g ...... 1 x 10^6 mg in a kg. So your conversion would be like this \[25kg \times \frac{ 1 \times 10^{6} mg}{ 1 kg }\]

OpenStudy (anonymous):

Oh wait .. I'm sorry. If you were to do it the long way .... you can convert kg to g the g to mg ... so the answer to your ? is 10^3 g /1kg \[25kg \times \frac{ 1 \times10^3g }{ 1kg }\times \frac{ 1 \times 10^3mg }{ 1g }\]